OpenStudy (hanner_b_nanner):

Will someone check my work that I have so far, and help me with the last one?? :D Fan&&Medal

OpenStudy (hanner_b_nanner):

Step 1: Calculations Main Show Tank Calculations: 1. The main tank has a radius of 70 feet. What is the volume of the quarter-sphered sized tank? Round your answer to the nearest whole number. You must explain your answer using words, and you must show all work and calculations to receive credit. The volume of the quarter-sphered tank rounds up to 478,676. You use the formula to find the volume, and that's how I found my answer. Volume of Sphere: V=4/3(3.14)R^3 V=4/3(3.14)(70)^3 V=4/3(1,077,020) V=478675.5555555556 Holding Tank Calculations 2. The holding tanks are congruent in size and both are in the shape of a cylinder that has been cut in half vertically. The bottom of the tank is a curved surface. What is the volume of both tanks if the radius of tank #1 is 15 feet and the height of tank #2 is 120 feet? You must explain your answer using words, and you must show all work and calculations to receive credit. The volume of both tanks is equal to 84,780. You use the formula, and plug in the numbers to get the answer. Volume of Cylinder: V=(3.14)R^2H V=(3.14)(15)^2(120) V=(3.14)(225)(120) V=84,780 Density Calculation: 3. In step 1, you found the volume (in cubic feet) of the main tank. If the maximum density of killer whales per cubic foot is 0.000011142, what is the maximum number of killer whales allowed in the main show tank at any given time? You must explain your answer using words, and you must show all work and calculations to receive credit. The maximum number of killer whales allowed in the main show tank would only be 5. You use the right formula, and plug in the numbers to get the right answer. D=mass/volume D=0.000011142 V=478675.5555555556 M=(0.000011142)(478675.5555555556) M= 5.33340304 ***4. The stadium is in need of a medical pool to attend to injured or sick whales. The pool will be located adjacent to the holding tanks. Your job is to design a medical pool using a 3-dimensional polyhedral shape you learned about in this module. For safety reasons, the tank can be no more than one-fourth the volume of a holding tank. You will need to submit a 2-dimensional drawing, the dimensions, the volume of your pool, and proof that the pool's volume is no more than one-fourth the volume of a holding tank to your instructor.

OpenStudy (anonymous):

The last calculation where you calculated the volume of the sphere is off, my calculator says it is 1436026.66667, and your number got smaller by multiplying by 4/3. Somehow xD and if it is quarter sized does that mean 1/4 of a sphere?

OpenStudy (hanner_b_nanner):

ohhh, ok! I'll try it again :D And yes, I think so?? Lol

OpenStudy (anonymous):

That means you should divide it by 4, or multiply by 1/4 that is the same thing. So that you find the volume of a quarter of the tank

OpenStudy (anonymous):

You were good all the way up to the last calculation on the first part so I wouldn't redo the whole thing

OpenStudy (hanner_b_nanner):

Alright! I'm just a little tired and my brain kinda stopped working :D Thank you!!!

OpenStudy (hanner_b_nanner):

ok!!!!

OpenStudy (anonymous):

Wait, ignore that I am wrong

OpenStudy (hanner_b_nanner):

ohh, alrighty :) Lol